# Smoke and Mirrors in Causal Downstream Tasks

### v1

@@ E.3 Results
Theory suggests that biased annotating criteria (i.e., depending on the experimental settings) can lead to biased treatment effect estimation, wrongly retrieving the conditional treatment effect on unseen experimental settings. Figure 8 validates this observation, and the results are validated via the $t$-tests reported in Table 7. Overall, the results perfectly align with the analogous discussion on [Anonimized]Ant .
@@
Distinct statistical and causal objectives cannot be used as a proxy for one another. We already formalized this in Lemma 3.1 and discussed it for [Anonimized]Ant dataset. In Figure 10 and 11, we systematically show it again for our new synthetic benchmark by comparing the rank-correlation among 200 ConvNets using random and biased sampling, respectively. Both matrices fully align with the discussion presented for the [Anonimized]Ant dataset in Section 6.

### v2

@@ E.3 Results
Theory suggests that biased annotating criteria (i.e., depending on the experimental settings) can lead to biased treatment effect estimation, wrongly retrieving the conditional treatment effect on unseen experimental settings. Figure 9 validates this observation, and the results are validated via the $t$-tests reported in Table 7. Overall, the results perfectly align with the analogous discussion on ISTAnt.
@@
Distinct statistical and causal objectives cannot be used as a proxy for one another. We already formalized this in Lemma 3.1 and discussed it for ISTAnt dataset. In Figure 11 and 12, we systematically show it again for our new synthetic benchmark by comparing the rank-correlation among 200 ConvNets using random and biased sampling, respectively. Both matrices fully align with the discussion presented for the ISTAnt dataset in Section 6.